# Primitive-Root Determinant Densities over Prime Fields and Implications for PRIM-LWE

### 11 Mar 2026 v1

@@ 5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
Then the values f(p1)f(p-1), as $p$ ranges over the primes, possess a continuous purely singular limiting distribution. More precisely, the empirical measures
@@
The shifted-prime analogue of the Erdős–Wintner theorem, due independently to Kátai [17] and Hildebrand [18], shows that a strongly additive function $h$ possesses a limiting distribution on the shifted primes p1p-1 if and only if the three classical Erdős–Wintner series
$\sum_{\begin{subarray}{c}\ell\ \textup{prime}\\ |h(\ell)|\leq 1\end{subarray}}\frac{h(\ell)}{\ell},\qquad\sum_{\begin{subarray}{c}\ell\ \textup{prime}\\ |h(\ell)|\leq 1\end{subarray}}\frac{h(\ell)^{2}}{\ell},\qquad\sum_{\begin{subarray}{c}\ell\ \textup{prime}\\ |h(\ell)|>1\end{subarray}}\frac{1}{\ell}$
converge; see Tenenbaum [2, pp. 257–258]. In the same reference it is noted that the shifted-prime limiting law is of pure type and is continuous if and only if
$\sum_{\begin{subarray}{c}\ell\ \textup{prime}\\ h(\ell)\neq 0\end{subarray}}\frac{1}{\ell}=\infty.$
We apply this to the strongly additive function
$f(n)=\sum_{\begin{subarray}{c}\ell\mid n\\ \ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
For every prime $\ell$,
$0<f(\ell)=\log\frac{\ell}{\ell-1}\leq\log 2<1,\qquad f(\ell)=\frac{1}{\ell}+O\!\left(\frac{1}{\ell^{2}}\right).$
Hence
$\sum_{\ell}\frac{f(\ell)}{\ell}<\infty,\qquad\sum_{\ell}\frac{f(\ell)^{2}}{\ell}<\infty,\qquad\sum_{|f(\ell)|>1}\frac{1}{\ell}=0.$
Therefore the values f(p1)f(p-1), as $p$ ranges over the primes, possess a limiting distribution $\mu_{f}$.
Since $f(\ell)\neq 0$ for every prime $\ell$ and $\sum_{\ell}1/\ell=\infty$, the limiting law $\mu_{f}$ is continuous.
Invoking the shifted-prime extension of Erdős’ singularity criterion (Tenenbaum [2, p. 258]; see also [3, p. 77]): for a strongly additive function $h$, once the limiting law of h(p1)h(p-1) exists, the condition
$\sum_{\ell\ \textup{prime}}\frac{h(\ell)^{2}}{\ell}<\infty$
forces that limiting law to be singular. Applying this with $h=f$, and using the already verified convergence of $\sum_{\ell}f(\ell)^{2}/\ell$, we conclude that $\mu_{f}$ is singular. Since $\mu_{f}$ is also continuous, it is purely singular. ∎
@@
By Theorem 7, the same is true for $c(p)$. Since the limiting law has support containing more than one point, it is not a point mass, and therefore $c(p)$ does not converge to a constant along the primes. ∎
@@ 6 Explicit Lower Bounds for Primes of Cryptographic Interest
The lower-bound argument is the same primorial/Mertens argument used in the proof of Corollary 5, but now with Theorem 8 replacing the crude bound $P_{n}(p)\geq\tfrac{1}{2}$.
Let
$m(x):=\max\{k\geq 1:N_{k}\leq x\}.$
Since $c(2)>0$, it suffices for large $x$ to consider odd primes $p\leq x$. For such $p$,
$N_{\omega(p-1)}\leq\prod_{r\mid(p-1)}r\leq p-1\leq x,$
so $\omega(p-1)\leq m(x)$. Therefore Theorem 8 gives
$c(p)\geq P_{3}\prod_{i=1}^{\omega(p-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\geq P_{3}\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr).$
As in the proof of Corollary 5, the relation
$N_{m(x)}\leq x<N_{m(x)+1}=N_{m(x)}\,p_{m(x)+1}$
implies $p_{m(x)}\sim\log x$. Hence, by Mertens’ theorem,
$\prod_{i=1}^{m(x)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\sim\frac{e^{-\gamma}}{\log p_{m(x)}}\asymp\frac{1}{\log\log x}.$
Therefore
$\min_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}c(p)\gg\frac{1}{\log\log x}.$
The matching upper bound is Proposition 1. ∎
The lower bound in Corollary 6 may be viewed as the prime-shifted specialization n=p1n=p-1 of the classical universal estimate
$\frac{\varphi(n)}{n}\gg\frac{1}{\log\log n},$
(see, e.g., Rosser and Schoenfeld [19]) together with the uniform bound $\prod_{j\geq 1}(1-p^{-j})\geq\tfrac{1}{2}$. The proof above keeps the primorial structure explicit, which makes it parallel to the upper-bound construction of Proposition 1.
@@
The comparison
$\min_{p\leq x}c(p)\asymp\frac{1}{\log\log x}$
should in principle be capable of being made fully explicit by combining explicit estimates for the Mertens product and for $\vartheta(x)$ (see, e.g., Rosser–Schoenfeld [19] and Dusart [20]) with an explicit least-prime-in-an-arithmetic-progression bound of Linnik type; see also Xylouris [21] for an admissible Linnik exponent. We do not carry out that bookkeeping here, since the present note only needs the qualitative asymptotic order.
@@
The bound $1/2$ for the infinite product can be sharpened for specific primes. For instance, $\prod_{j\geq 1}(1-q^{-j})\geq 1-1/(q-1)$, and this lower bound exceeds $2/3$ for $q\geq 5$ and is at least $0.99$ for $q\geq 101$. For large implementation-oriented NTT primes (for example $q>2^{30}$), this factor is numerically indistinguishable from $1$. The current NIST-standardized lattice moduli $q=3329$ and $q=8380417$ are much smaller [4, 5], but even there the same elementary bound already gives $\prod_{j\geq 1}(1-q^{-j})>0.9996$ and $\prod_{j\geq 1}(1-q^{-j})>0.9999998$, respectively.
@@
If $\omega(q-1)=1$ and $q$ is an odd prime, then q1q-1 is an even prime power, hence $q-1=2^{m}$ for some $m\geq 1$; equivalently, $q$ is a Fermat prime. In this case
@@
By Dusart’s explicit product estimate [20, Theorem 5.9], which we use conservatively with threshold $x\geq 2\,278\,383$,
@@
Since $q>2^{30}$, we have q1>2 278 383q-1>2\,278\,383 and $\log(q-1)>20$, so the parenthetical factor lies in $(1-2.5\times 10^{-5},\;1+2.5\times 10^{-5})$.
@@
The bound $c(q)>0.56/\!\log q$ of Corollary 8 is a simple, fully explicit estimate obtained by taking the Mertens product over all primes up to q1q-1. However, the sharper universal pointwise bound
$c(q)\;\geq\;P_{3}\prod_{i=1}^{\omega(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\geq\;P_{3}\prod_{i=1}^{m(q-1)}\Bigl(1-\frac{1}{p_{i}}\Bigr)\;\gg\;\frac{1}{\log\log q}$
follows directly from Theorem 8 together with $\omega(q-1)\leq m(q-1)$ and Mertens’ theorem, where $m(n):=\max\{k:N_{k}\leq n\}$. This matches the sharp order established in Corollary 6 and shows that $1/\!\log\log q$, not $1/\!\log q$, is the correct worst-case scale for the rejection-sampling overhead.
@@
Let $\Phi$ denote the standard normal distribution function. The shifted-prime analogue of the Erdős–Kac theorem (due to Barban [22]; see also Halberstam [23] and Elliott [24, Ch. 12]) asserts that for every real $z$,
@@
For the two prime moduli appearing, across all parameter sets, in the current NIST standards for ML-KEM and ML-DSA [4, 5], Table 1 records both the general $\omega(q-1)$-based bounds and the sharper exact-factorization bounds. The fourth column records the density lower bound $B(\omega(q-1))$, so the corresponding general rejection-sampling overhead upper bound is its reciprocal $1/B(\omega(q-1))$; for the two rows of Table 1 these are $1/B(2)\approx 5.36$ and $1/B(4)\approx 7.82$, respectively. The exact factorizations give
$\frac{\varphi(3328)}{3328}=\frac{6}{13},\qquad\frac{\varphi(8380416)}{8380416}=\frac{100}{341},$
so together with $P(q)\geq 1-\frac{1}{q-1}$ one obtains
@@
Direct evaluation of the defining product shows that these lower bounds are already essentially exact:
$c(3329)=\frac{6}{13}\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{3329^{j}}\Bigr)\approx 0.46139978,\qquad c(8380417)=\frac{100}{341}\prod_{j=1}^{\infty}\Bigl(1-\frac{1}{8380417^{j}}\Bigr)\approx 0.29325510.$
Equivalently, the corresponding dimension-uniform expected rejection-sampling overheads $1/c(q)$ are approximately $2.1674$ and $3.4101$. For each fixed dimension $n$, the exact overheads $1/c_{n}(q)$ are no larger.
@@
Consider $q=2^{64}-2^{32}+1=18\,446\,744\,069\,414\,584\,321$, a well-known NTT-friendly prime. This modulus is not the native modulus of the standardized NIST-PQC schemes, but it is used as a 64-bit NTT-friendly ambient field in implementation-oriented work on lattice cryptography, so it is a natural cryptographic test case [25]. One verifies that $q-1=2^{32}\cdot 3\cdot 5\cdot 17\cdot 257\cdot 65537$, so $\omega(q-1)=6$ and
$\frac{\varphi(q-1)}{q-1}=\frac{1}{2}\cdot\frac{2}{3}\cdot\frac{4}{5}\cdot\frac{16}{17}\cdot\frac{256}{257}\cdot\frac{65536}{65537}\approx 0.2500.$
Since $P(q)>1-10^{-19}$ for such large $q$, the limiting density $c(q)$ is about $0.25$. The exact finite-$n$ density $c_{n}(q)$ is then also essentially $0.25$, so about one in four random $n\times n$ matrices has a primitive-root determinant, and the security reduction loses a factor of about $4$.

### 18 Mar 2026 v2

@@ 5.2 From $\varphi(p-1)/(p-1)$ to $c(p)$
Lemma 6 (Shifted-prime pure-type and continuity criterion).
Let $f$ be a real strongly additive arithmetical function satisfying the three convergence conditions of the shifted-prime Erdős–Wintner theorem, so that f(p1)f(p-1) possesses a limiting distribution $\mu_{f}$ over the primes. Then $\mu_{f}$ is of pure type: it is either purely discrete, purely singular, or absolutely continuous with respect to Lebesgue measure. Moreover, $\mu_{f}$ is continuous (i.e. has no atoms) if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$.
Proof.
The relevant shifted-prime Erdős–Wintner framework is due to Kátai [14] and Hildebrand [15] for the shift $a=1$; the methods extend to any fixed nonzero shift via the Bombieri–Vinogradov theorem, and for the shift a=1a=-1 used here, the explicit characteristic function is recorded in Tenenbaum [3, formula (3), p. 258]. The same conclusions hold for general (not necessarily strongly) additive functions under the Erdős–Wintner hypotheses, by the identical Jessen–Wintner argument applied to the full local factors; we state the strongly additive case since it is all that we require.
By Tenenbaum [3, formula (3), p. 258], the characteristic function of $\mu_{f}$ is given by
$\widehat{\mu_{f}}(\tau)\;=\;\prod_{p}\!\left(1-\frac{1}{p-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(p^{\nu})}}{p^{\nu}}\right)$
The shift is p1p-1, corresponding to a=1a=-1 in the notation of [3, formula (3)]. Since no prime divides 1-1, the product runs over all primes. For the shift a=1a=-1, the local factor at $\ell=2$ in the characteristic function reduces, when $f$ is strongly additive, to $1-\frac{1}{2-1}+\sum_{\nu\geq 1}\frac{e^{i\tau f(2)}}{2^{\nu}}=e^{i\tau f(2)}$, which is a deterministic translation by $f(2)$. This factor carries no randomness and does not affect pure type or continuity; it merely shifts the law. It therefore suffices to apply the Jessen–Wintner theorem [16] (see also Tenenbaum [13, Theorem 2.7(b)]) to the infinite convolution over the odd primes. Each odd-prime factor is the characteristic function of a probability measure supported on a countable set, so the Jessen–Wintner theorem gives that the resulting infinite convolution is of pure type. The continuity criterion follows from the same theorem: the convolution is continuous if and only if $\sum(1-\sigma_{\ell})=\infty$, where $\sigma_{\ell}$ is the maximal atom of the factor at $\ell$. Since $f$ is strongly additive, the local factor at an odd prime $\ell$ places mass $(\ell-2)/(\ell-1)$ at the origin and mass $1/(\ell-1)$ at $f(\ell)$. When $f(\ell)\neq 0$, these two points are distinct, so $\sigma_{\ell}=(\ell-2)/(\ell-1)$ and $1-\sigma_{\ell}=1/(\ell-1)\asymp 1/\ell$. It follows that  $\sum_{\ell\text{ odd}}(1-\sigma_{\ell})$ diverges if and only if $\sum_{f(\ell)\neq 0}1/\ell=\infty$, since the prime $\ell=2$ does not enter the odd-prime convolution. Since a deterministic translation preserves pure type and continuity, the same conclusions hold for $\mu_{f}$. ∎
Lemma 7 (Prime density for a prescribed smooth kernel).
Let $y\geq 3$, and for each prime $p$ define $a_{y}(p-1):=\prod_{\begin{subarray}{c}\ell\leq y,\;\ell\textup{ prime}\\ \ell\mid(p-1)\end{subarray}}\ell$. Let $m$ be an even squarefree positive integer with $P^{+}(m)\leq y$. Then the set $\{p\textup{ prime}:a_{y}(p-1)=m\}$ has relative density
$\delta_{y}(m)\;=\;\frac{1}{\varphi(m)}\prod_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\nmid m\end{subarray}}\frac{\ell-2}{\ell-1}$
among the primes, in the sense that
$\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:p\textup{ prime},\;a_{y}(p-1)=m\bigr\}\;=\;\delta_{y}(m).$
Moreover, writing $C_{y}:=\prod_{3\leq\ell\leq y}(\ell-2)/(\ell-1)\asymp 1/\!\log y$, we have $\delta_{y}(m)\geq C_{y}/\varphi(m)$.
Proof.
For each odd prime $\ell\leq y$ with $\ell\mid m$, the condition $\ell\mid(p-1)$ forces $p\equiv 1\pmod{\ell}$. This selects one residue class out of $\ell-1$ reduced classes modulo $\ell$. For each odd prime $\ell\leq y$ with $\ell\nmid m$, the condition $\ell\nmid(p-1)$ excludes the class $p\equiv 1\pmod{\ell}$, leaving $\ell-2$ admissible classes out of $\ell-1$. The prime $\ell=2$ imposes no constraint since $2\mid(p-1)$ for every odd prime $p$ and $2\mid m$ by hypothesis. By the Chinese Remainder Theorem, these local conditions combine to specify a union of reduced residue classes modulo the primorial $M_{y}:=\prod_{\ell\leq y,\;\ell\text{ prime}}\ell$. The finitely many primes $p\leq y$ need not satisfy these residue-class conditions, but they contribute $O_{y}(1)$ to the count and do not affect the relative density. Since $M_{y}$ is fixed for each $y$, the Siegel–Walfisz theorem (Theorem 5) gives the stated density.
Since each $m$ is even and squarefree, $\varphi(m)=\prod_{\begin{subarray}{c}\ell\mid m\\ \ell>2\end{subarray}}(\ell-1)$, so the first product in the expression for $\delta_{y}(m)$ equals $1/\varphi(m)$. The lower bound follows because each factor $(\ell-2)/(\ell-1)\in(0,1)$. To justify the asymptotic $C_{y}\asymp 1/\!\log y$, write
$\log C_{y}=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\log\frac{\ell-1}{\ell-2}=\sum_{\begin{subarray}{c}3\leq\ell\leq y\\ \ell\text{ prime}\end{subarray}}\frac{1}{\ell}+O(1),$
since $\log((\ell-1)/(\ell-2))=1/\ell+O(1/\ell^{2})$ and $\sum_{\ell}1/\ell^{2}<\infty$. By Mertens’ second theorem the sum over primes equals $\log\log y+O(1)$, so $C_{y}=e^{-\log\log y+O(1)}\asymp 1/\!\log y$. ∎
Lemma 8 (Tail-mean estimate for f(p1)f(p-1)).
Let $f(n):=\log(n/\varphi(n))$, and for each prime $p$ and parameter $y\geq 3$ define
$T_{y}(p):=\sum_{\begin{subarray}{c}\ell\mid(p-1)\\ \ell>y,\;\ell\textup{ prime}\end{subarray}}\log\frac{\ell}{\ell-1}.$
Then, for each fixed $y\geq 3$,
$\lim_{x\to\infty}\frac{1}{\pi(x)}\sum_{\begin{subarray}{c}p\leq x\\ p\textup{ prime}\end{subarray}}T_{y}(p)\;=\;\eta(y),$
where $\eta(y):=\sum_{\ell>y,\;\ell\textup{ prime}}f(\ell)/(\ell-1)\ll 1/y$. In particular, $\eta(y)\to 0$ as $y\to\infty$.
Proof.
Since $f(\ell)=1/\ell+O(1/\ell^{2})$, we have $\eta(y)=\sum_{\ell>y}1/(\ell(\ell-1))+O(\sum_{\ell>y}1/\ell^{3})\ll 1/y$. Interchanging sums gives
$\sum_{p\leq x}T_{y}(p)\;=\;\sum_{\begin{subarray}{c}y<\ell<x\\ \ell\text{ prime}\end{subarray}}f(\ell)\,\pi(x;\ell,1).$
Fix $A:=2$ and split the sum at the threshold $L:=(\log x)^{A}$.
Small primes ($y<\ell\leq L$). Each such $\ell$ satisfies $\ell\leq(\log x)^{A}$, so by the Siegel–Walfisz theorem (Theorem 5),
$\pi(x;\ell,1)\;=\;\frac{\operatorname{Li}(x)}{\ell-1}\;+\;E(x;\ell,1),\qquad|E(x;\ell,1)|\;\ll_{A}\;xe^{-c_{A}\sqrt{\log x}}.$
Since $f(\ell)\leq 1$ for every prime $\ell$ and $\{\text{primes }\ell\in(y,L]\}\leq L$, the total error is
$\sum_{y<\ell\leq L}f(\ell)\,|E(x;\ell,1)|\;\leq\;L\cdot O_{A}\!\bigl(xe^{-c_{A}\sqrt{\log x}}\bigr)\;=\;o\bigl(\pi(x)\bigr).$
The main term from this range is $\operatorname{Li}(x)\sum_{y<\ell\leq L}f(\ell)/(\ell-1)=\operatorname{Li}(x)\bigl(\eta(y)+O(1/L)\bigr)$, since the tail $\sum_{\ell>L}f(\ell)/(\ell-1)\ll 1/L$.
Large primes ($\ell>L$). For every prime $\ell\leq x$, the trivial bound $\pi(x;\ell,1)\leq x/\ell+1$ and $f(\ell)=\log(\ell/(\ell-1))\leq 2/\ell$ give
$\sum_{\ell>L}f(\ell)\,\pi(x;\ell,1)$ $\;\leq\;\sum_{L<\ell\leq x}\frac{2}{\ell}\left(\frac{x}{\ell}+1\right)\;\leq\;2x\sum_{\ell>L}\frac{1}{\ell^{2}}\;+\;2\sum_{\ell\leq x}\frac{1}{\ell}\;\ll\;\frac{x}{L}\;+\;\log\log x.$
Dividing by $\pi(x)\sim x/\!\log x$, this contributes $O(\log x/L)+o(1)=o(1)$. Combining both ranges and dividing by $\pi(x)$ yields the claim. ∎
Lemma 9 (Shifted-prime pure singularity for $\log(n/\varphi(n))$).
Let
The shifted-prime limiting measure $\mu_{f}$ is purely singular with respect to Lebesgue measure. Here $\mu_{f}$ denotes the weak limit of the empirical measures $\pi(x)^{-1}\sum_{p\leq x}\delta_{f(p-1)}$.
Proof.
We adapt the integer-case argument of Erdős [5] (see also [13, Exercise 256] with solution in [17, Solution 2.4, pp. 147–148]) to the shifted-prime setting, using the Siegel–Walfisz theorem (Theorem 5) in place of the integer equidistribution input.
By the shifted-prime Erdős–Wintner theorem (Tenenbaum [3, pp. 257–258] for the shift a=1a=-1; Kátai [14], Hildebrand [15] for the original $a=1$ case), the limiting measure $\mu_{f}$ exists. Indeed, by Lemma 3 the three convergence hypotheses are satisfied and the continuity criterion holds, so the limiting measure $\mu_{f}$ exists and is continuous. By Lemma 6, $\mu_{f}$ is of pure type. Since it is continuous (as verified above), it is therefore either purely singular or absolutely continuous. We show that $\mu_{f}$ is not absolutely continuous, which forces pure singularity.
Step 1: Tail estimate. Throughout the proof, the single prime $p=2$ is negligible in any asymptotic density statement. Fix a parameter $\delta\in(0,\tfrac{1}{2})$ and, for $y\geq 3$, set $R:=\lfloor y^{\delta}\rfloor$. Since $f$ is strongly additive, for each prime $p$ we decompose $f(p-1)=f(a_{y}(p-1))+T_{y}(p)$, where $a_{y}(p-1)$ and $T_{y}(p)$ are as in Lemma 8. By that lemma and Markov’s inequality,
$\limsup_{x\to\infty}\;\frac{1}{\pi(x)}\#\bigl\{p\leq x:T_{y}(p)>1/R^{2}\bigr\}\;\leq\;R^{2}\,\eta(y)\;\ll\;y^{2\delta-1},$ (1)
which tends to 0 as $y\to\infty$ (since $\delta<\tfrac{1}{2}$). Here and in Steps 2–3, the parameter $y$ is held fixed while $x\to\infty$; the conclusions are then strengthened by letting $y\to\infty$.
Step 2: A set of small Lebesgue measure carrying positive relative prime density. Let $m_{1}<m_{2}<\cdots<m_{R}$ be the first $R=\lfloor y^{\delta}\rfloor$ even squarefree positive integers. Since the counting function of even squarefree integers satisfies $A(t):=\#\{m\leq t:m\text{ even squarefree}\}\sim(2/\pi^{2})\,t$, for any fixed $C>\pi^{2}/2$ we have $A(Cj)\geq j$ for all sufficiently large $j$. After enlarging $C$ if necessary, we may assume $m_{j}\leq Cj$ for all $j\geq 1$. In particular $m_{R}\leq CR$. Therefore $P^{+}(m_{j})\leq m_{j}\leq Cy^{\delta}<y$ for all sufficiently large $y$ (since $\delta<1$). Moreover,
$\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\geq\;\sum_{j=1}^{R}\frac{1}{m_{j}}\;\geq\;\frac{1}{C}\sum_{j=1}^{R}\frac{1}{j}\;\gg\;\log R\;=\;\delta\log y+O(1).$ (2)
By Lemma 7, the relative density of $\{p:a_{y}(p-1)=m_{j}\}$ among the primes is $\delta_{y}(m_{j})\geq C_{y}/\varphi(m_{j})$, where $C_{y}\asymp 1/\!\log y$. Combining with (2),
$\sum_{j=1}^{R}\delta_{y}(m_{j})\;\geq\;C_{y}\sum_{j=1}^{R}\frac{1}{\varphi(m_{j})}\;\gg\;\frac{\delta\log y}{\log y}\;=\;\delta\;>\;0.$ (3)
Since the $m_{j}$ are distinct, the sets $\{p:a_{y}(p-1)=m_{j}\}$ are pairwise disjoint, so the relative density of their union $\mathcal{A}_{y}:=\{p:a_{y}(p-1)\in\{m_{1},\ldots,m_{R}\}\}$ is the sum $\sum_{j}\delta_{y}(m_{j})$. In particular, there exists an absolute constant $c_{0}>0$ (depending only on $\delta$) such that for all sufficiently large $y$, the relative density of $\mathcal{A}_{y}$ among the primes is at least $c_{0}$.
Define the closed set
$\bar{I}_{y}\;:=\;\bigcup_{j=1}^{R}\bigl[\,f(m_{j}),\;f(m_{j})+1/R^{2}\,\bigr].$
By subadditivity of Lebesgue measure, $\lambda(\bar{I}_{y})\leq R\cdot(1/R^{2})=1/R=1/\lfloor y^{\delta}\rfloor\to 0$ as $y\to\infty$. For any prime $p$ with $a_{y}(p-1)=m_{j}$ and $T_{y}(p)\leq 1/R^{2}$, the identity $f(p-1)=f(a_{y}(p-1))+T_{y}(p)$ and the nonnegativity of $T_{y}(p)$ give $f(p-1)\in[f(m_{j}),\,f(m_{j})+1/R^{2}]\subset\bar{I}_{y}$.
Step 3: Contradiction with absolute continuity. Recall that $y$ (and hence $R$) are still fixed. By (1),
$\limsup_{x\to\infty}\;\frac{1}{\pi(x)}\#\bigl\{p\leq x:T_{y}(p)>1/R^{2}\bigr\}\;=\;o_{y}(1).$
Hence, for all sufficiently large $y$,
$\liminf_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:f(p-1)\in\bar{I}_{y}\bigr\}\;\geq\;c_{0}-o_{y}(1)\;\geq\;\frac{c_{0}}{2}\;>\;0.$
[…]
$\mu_{f}(\bar{I}_{y})\;=\;\lim_{x\to\infty}\frac{1}{\pi(x)}\#\bigl\{p\leq x:f(p-1)\in\bar{I}_{y}\bigr\}\;\geq\;\frac{c_{0}}{2}\;>\;0.$
If $\mu_{f}$ were absolutely continuous with density $h\in L^{1}(\mathbb{R})$, then by the absolute continuity of the Lebesgue integral, $\mu_{f}(\bar{I}_{y})=\int_{\bar{I}_{y}}h\,\mathrm{d}t\to 0$ as $\lambda(\bar{I}_{y})\to 0$, contradicting $\mu_{f}(\bar{I}_{y})\geq c_{0}/2$ for all large $y$. Therefore $\mu_{f}$ is not absolutely continuous, and being of pure type and continuous, it is purely singular. ∎
@@
Then the empirical measures
@@
By Lemma 3, the three convergence conditions of the shifted-prime Erdős–Wintner theorem (Kátai [14], Hildebrand [15] for $p+1$; Tenenbaum [3, pp. 257–258] for the general shift including p1p-1) are satisfied, and the continuity criterion holds. Therefore f(p1)f(p-1) possesses a continuous limiting distribution $\mu_{f}$, which is of pure type by Lemma 6.
It remains to rule out absolute continuity. By Lemma 9, $\mu_{f}$ is purely singular. ∎
@@
By Theorem 7, the same is true for $c(p)$. ∎
@@ 6 Explicit Lower Bounds for Primes of Cryptographic Interest
The lower bound follows by the same primorial/Mertens argument used in the proof of Corollary 5: for every odd prime $p\leq x$, $\omega(p-1)\leq m(x)$, and Lemma 1 together with Mertens’ theorem gives $\min_{p\leq x,\,p\textup{ prime}}c(p)\gg 1/\!\log\log x$. Theorem 8 sharpens the constant from $\tfrac{1}{2}$ to $P_{3}$ but does not affect the asymptotic order. The matching upper bound is Proposition 1. ∎
@@
The asymptotic $\min_{p\leq x}c(p)\asymp 1/\!\log\log x$ should be capable of being made fully explicit by combining explicit Mertens product and $\vartheta(x)$ estimates (e.g., Rosser–Schoenfeld [18], Dusart [19]) with an explicit Linnik-type bound (see Xylouris [20]). We do not pursue the bookkeeping here.
@@
The bound $1/2$ for the infinite product can be sharpened for specific primes: $\prod_{j\geq 1}(1-q^{-j})\geq 1-1/(q-1)$, which exceeds $2/3$ for $q\geq 5$ and is at least $0.99$ for $q\geq 101$. For large NTT primes with $q>2^{30}$, this factor is numerically indistinguishable from 11. Even for the smaller NIST moduli $q=3329$ and $q=8380417$ [21, 22], the elementary bound gives $\prod_{j\geq 1}(1-q^{-j})>0.9996$ and $\prod_{j\geq 1}(1-q^{-j})>0.9999998$, respectively.
@@
If $\omega(q-1)=1$ and $q$ is an odd prime, then q1q-1 is an even prime power, hence $q-1=2^{m}$ for some $m\geq 1$. Equivalently, $q$ is a Fermat prime. In this case
@@
By Dusart’s explicit product estimate [19, Theorem 5.9], which we use conservatively with threshold $x\geq 2\,278\,383$,
@@
Since $q>2^{30}$, we have q1>2 278 383q-1>2\,278\,383 and $\log(q-1)>20$, so the correction factor lies in $(1-2.5\times 10^{-5},\;1+2.5\times 10^{-5})$.
@@
The bound $c(q)>0.56/\!\log q$ of Corollary 8 is a simple, fully explicit estimate obtained by taking the Mertens product over all primes up to q1q-1. The sharper universal pointwise bound $c(q)\gg 1/\!\log\log q$ of Corollary 7 shows that $1/\!\log\log q$, not $1/\!\log q$, is the correct worst-case scale for the rejection-sampling overhead.
@@
Let $\Phi$ denote the standard normal distribution function. The shifted-prime analogue of the Erdős–Kac theorem (due to Barban [23]; see also Halberstam [24] and Elliott [25, Ch. 12]) asserts that for every real $z$,
@@
For the two prime moduli appearing, across all parameter sets, in the current NIST standards for ML-KEM and ML-DSA [21, 22], Table 1 records both the general $\omega(q-1)$-based bounds and the sharper exact-factorization bounds. The exact factorizations give $\varphi(3328)/3328=6/13$ and $\varphi(8380416)/8380416=100/341$. Together with $P(q)\geq 1-\frac{1}{q-1}$, one obtains
@@
A short tail-bound argument (taking $J=2$ partial products of $P(q)$ and using the estimate $P_{J}(q)-P(q)\leq q^{-J}/(q-1)$) confirms that these lower bounds are tight to many decimal places:
$c(3329)=0.4613997781\ldots,\qquad c(8380417)=0.2932550969\ldots.$
The corresponding dimension-uniform expected rejection-sampling overheads $1/c(q)$ are approximately $2.17$ and $3.41$. For each fixed dimension $n$, the exact overheads $1/c_{n}(q)$ are no larger.
Context $q$ factorization of q1q-1 $\omega(q-1)$ $1/c(q)$
ML-KEM (FIPS 203) $3329$ $^{8}\cdot 13 22$ $2.17$
ML-DSA (FIPS 204) $8380417$ $2^{13}\cdot 3\cdot 11\cdot 31$ 44 $3.41$
NewHope / FALCON [^1] FALCON is being standardized by NIST as FN-DSA (FIPS 206). [26, 27] $12289$ $^{12}\cdot 3 22$ $3.00$
BFV/BGV FHE (small) [28, 29, 30] $786433$ $^{18}\cdot 3 22$ $3.00$
SEAL BFV default [31] $132120577$ $2^{21}\cdot 3^{2}\cdot 7$ 33 $3.50$
BFV/BGV FHE (medium) [28, 29, 30] $2013265921$ $2^{27}\cdot 3\cdot 5$ 33 $3.75$
NTTRU / BLISS [32, 33] $7681$ $2^{9}\cdot 3\cdot 5$ 33 $3.75$
64-bit NTT prime [34] $2^{64}-2^{32}+1$ $2^{32}\cdot 3\cdot 5\cdot 17\cdot 257\cdot 65537$ 66 $4.00$
Fermat prime ($2^{16}+1$) $65537$ $2^{16}$ 11 $2.00$
Table 2: Representative NTT-friendly primes from lattice-based cryptography and FHE implementations, together with $\omega(q-1)$ and the dimension-uniform expected rejection-sampling overhead $1/c(q)$. All overheads are at most $4.00$. The first two rows reproduce the NIST entries from Table 1. The SEAL entry is the default coefficient modulus for poly_modulus_degree = 1024 at the 128-bit security level.
Table 2 collects nine representative moduli drawn from current lattice-based and FHE implementations. Every modulus in the table satisfies $\omega(q-1)\leq 6$ and incurs a dimension-uniform rejection-sampling overhead $1/c(q)\leq 4.00$. The table substantiates the claim that the primes typically encountered in practice have controlled factorization structure, which yields a good lower bound on $c(q)$.
@@
For the 64-bit NTT prime $q=2^{64}-2^{32}+1$ (sometimes called the “Goldilocks prime” in the zero-knowledge community, not to be confused with the Ed448-Goldilocks prime $2^{448}-2^{224}-1$ [35]), used as an NTT-friendly ambient field in implementation-oriented lattice cryptography [34], one has $\omega(q-1)=6$ and $\varphi(q-1)/(q-1)\approx 0.2500$. Since $P(q)>1-10^{-19}$, one obtains $c(q)\approx 0.25$, so roughly a quarter of random $n\times n$ matrices have a primitive-root determinant and the security reduction loses a factor of about $4$.